OpenStudy (anonymous):

solve the problems. n+9=27 solve for letter.

OpenStudy (anonymous):

n=18

OpenStudy (ash2326):

we've to find n n+9=27 n=27-9 n=18

OpenStudy (anonymous):

thanks here is another one in a mintue.

OpenStudy (anonymous):

n=18